Circular No. 7239

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA IN THE SMALL MAGELLANIC CLOUD J.-F. Glicenstein (DAPNIA, Saclay), on behalf of the EROS2 collabo-

ration, reports the discovery of an apparent nova in the Small Magellanic Cloud. The object, located at α = 00^{h}53^{m}40^{s}.0, δ = −72^{o}12′32′′ (equinox 2000.0), triggered the EROS2 microlensing alert system and peaked in brightness sometime between consecutive observations on 1999 July 26.1 UT (fainter than V ∼ 20.6) and 28.1 UT (R_{c} = 13.6). The object had faded to R_{c} = 15.6 by Aug. 13 UT. In the pre-outburst field, the closest stellar object (R_{c} = 20.2) was some 5′′ distant from the location of the nova. Find-

ing charts can be retrieved from http://www-dapnia.cea.fr/Phys/Spp/
Experiences/EROS/novae/index.html.

SUPERNOVAE 1999dg AND 1999dh A. V. Filippenko, University of California at Berkeley, reports that M.

R. Metzger and T. A. Small (Caltech) obtained CCD spectra (range 390– 880 nm, resolution 1 nm) of SNe 1999dg and 1999dh (cf. IAUC 7229) on Aug. 9 UT with the Keck-II 10-m telescope. Inspection of the uncalibrated data reveals that SN 1999dg is of type Ia, within about a week past maximum brightness. SN 1999dh is a normal type II SN, perhaps a month or two past maximum, in agreement with the report of Salvo et al. (IAUC 7238).

SUPERNOVAE 1999df, 1999di AND 1999dj A. V. Filippenko, T. Matheson, R. T. Chornock and A. L. Coil, Univer-

sity of California at Berkeley, report that uncalibrated CCD spectra (range 330–1000 nm) obtained on Aug. 17 UT with the Shane 3-m reflector at Lick Observatory reveal that SN 1999df (cf. IAUC 7228) is of type II. Broad Hα emission is easily visible in the noisy data. The approximate redshift is 0.038, based on narrow emission lines in the host galaxy. SN 1999di (cf. IAUC 7234) is of type Ib, probably within a few weeks past maximum, with prominent He I absorption lines at a velocity of 8100 km/s. The identity of relatively strong absorption at rest wavelength 630 nm is uncertain; it might be C II 658 nm, Si II 635 nm, or possibly even Hα 656 nm. If the absorption is due to Hα (at velocity 12 000 km/s), then the object would be of type IIb. The redshift is 0.016, again from narrow lines in the host. SN 1999dj (cf. IAUC 7234) is of type II, with well-developed P-Cyg profiles. The redshift is 0.015.

1999 August 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Gareth V. Williams